Suppose that a bond pays semiannual coupons in the sequence 35, 40, 35, 45, 35, 50, ... , 35, 95. If the redemption value is 2500 dollars and the yield rate is 8.2 percent convertible semiannually, what is the price of the bond?
Answer #1

Interest rate for 5 months = 8.2%/2 = 4.1%

The bond pays alternating payments of 35 and an increasing AP starting from 40

No. of periods of increasing AP= (95-40)/5 + 1 = 11+1 = 12

and there are 12 payments of 35

So, total period = 24 or 12 years

So,

P = 35/1.041 + 40/1.041^2+ .... + 95/1.041^24 + 2500/1.041^24

= $ 1686.7395 (do it in excel sheet)